# The Bisognano-Wichmann property for non-unitary Wightman conformal field theories

### v1

@@ 3 Analytic continuation for non-unitary Wightman CFTs
We have thus shown that for $\tau\in{\mathbb{S}}$ the domain $D(\widetilde{V}_{\tau})$ contains ${\mathcal{P}}_{\mathcal{X}}(I_{+})\Omega$, where we recall ${\mathcal{P}}_{\mathcal{X}}(I_{+})\subset{\mathcal{P}}(I_{-})$ is generated by operators $\varphi(F|_{I_{+}})$ with $F\in{\mathcal{X}}$. We would like to extend this result to ${\mathcal{P}}(I_{+})\Omega$ via a limiting argument, approximating $f\in C^{\infty}_{0}(I_{+})$ by functions of the form $F|_{I_{+}}$. We will begin by examining the growth rate of the function $\lambda(V_{t}\Phi)$.
@@
    where $\left\|f\right\|_{C^{N}}=\sum_{j=1}^{N}\big{\|}\tfrac{d^{j}}{d\theta^{j}}f\big{\|}_{\infty}$.
@@
is jointly continuous on $C^{\infty}(S^{1})^{k}$. Thus by standard results (c.f. the beginning of [Trè67, §41]) there exists a positive integer $N$ such that
@@
where we take the convention that the (non-commuting) products are ordered left-to-right. Hence
@@
$\qquad\leq\sum_{j=1}^{k}\left|\lambda\left(\left[\prod_{\ell=1}^{j-1}\varphi_{\ell}(\beta_{d_{j}}(v_{t})f_{\ell})\right]\varphi_{j}(\beta_{d_{j}}(v_{t})[f_{j}-g_{j}])\left[\prod_{\ell=j+1}^{k}\varphi_{\ell}(\beta_{d_{j}}(v_{t})g_{\ell})\right]\Omega\right)\right|\leq$
@@
by (3.6), and similarly for the $g_{j}$. By Lemma 3.8 applied to the functions $f_{j,s}$ and $g_{j,s}$ (along with the fact that $\widetilde{V}_{t+is}=V_{t}\widetilde{V}_{is}$) we have
@@
for some positive integer $N$ and some function $C$ with at most exponential growth that does not depend on the functions $F_{j}$ and $G_{j}$. Since $F$ and $G$ are smooth on ${\mathbb{O}}$, the maximum
@@
is achieved. Hence there is a constant $K>0$ (depending on the $F_{j}$ and $G_{j}$) such that
@@
where $M=\max_{t}e^{-t^{2}}C(t)$ does not depend on the functions $F_{j}$,$G_{j}$.
@@
$\big{|}\lambda\big{(}\big{[}\widetilde{V}_{\tau}$ $\varphi_{1}(f_{1,n})\cdots\varphi_{k}(f_{k,n})\Omega\big{]}-\big{[}V_{\tau}\varphi_{1}(f_{1,m})\cdots\varphi_{k}(f_{k,m})\Omega\big{]}\big{)}\big{|}$
@@
$G_{\lambda,n}(\tau)=\lambda\big{(}V_{\tau}\varphi_{1}(f_{1,n})\cdots\varphi_{k}(f_{k,n})\Omega\big{)}$
@@ 4 Invariant Hermitian forms, Bisognano-Wichmann, and the KMS condition
[…]
@@
and moreover we have $(-1)^{d_{\varphi}}\varphi^{\dagger}\in{\mathcal{F}}$. We also have
$\langle U(\gamma)\Phi,U(\gamma)\Psi\rangle=\langle\Phi,\Psi\rangle,\qquad\Phi,\Psi\in{\mathcal{D}},\,\gamma\in\operatorname{M\ddot{o}b}.$
The sesquilinear form is assumed to be jointly ${\mathcal{F}}$-strong continuous and is automatically Hermitian symmetric, meaning $\langle\Phi,\Psi\rangle=\overline{\langle\Psi,\Phi\rangle}$ for all $\Phi,\Psi\in{\mathcal{D}}$. Given $\theta$, the existence of a unique sesquilinear form on ${\mathcal{D}}$ with these properties is equivalent to $\langle\,\cdot\,,\theta\,\cdot\,\rangle$ being a nondegenerate invariant bilinear form on the Möbius vertex algebra underlying ${\mathcal{F}}$. The theory is called unitary if the Hermitian form is an inner product (normalized so that $\left\|\Omega\right\|=1$). In terms of the correspondence between Wightman CFTs and vertex algebras, an involutive Wightman CFT corresponds to a Möbius vertex algebra with a sesquilinear form and antiunitary operator $\theta$ such that $\langle\,\cdot\,\theta\,\cdot\,\rangle$ is an invariant bilinear form in the sense of [FHL93, Li94]. See [CRTT24, Def. 4.9] and the surrounding discussion for more detail.
Given a continuous map $x\in\operatorname{End}({\mathcal{D}})$, we say that $x$ has an adjoint if there exists $x^{*}\in\operatorname{End}({\mathcal{D}})$ such that $\langle x\Phi,\Psi\rangle=\langle\Phi,x^{*}\Psi\rangle$ for all $\Phi,\Psi\in{\mathcal{D}}$ [^8] See also the discussion following Corollary 4.3 regarding the partially defined adjoint of a partially defined (and not necessarily continuous) linear map., and write $\operatorname{End}_{*}({\mathcal{D}})$ for the collection of continuous operators $x\in\operatorname{End}({\mathcal{D}})$ that have an adjoint $x^{*}\in\operatorname{End}({\mathcal{D}})$. In light of (4.1) we see that $\varphi(f)\in\operatorname{End}_{*}({\mathcal{D}})$, and moreover the algebras ${\mathcal{P}}(I)$ are *-algebras, i.e., they are closed under taking adjoints. We have $\theta{\mathcal{P}}(I)\theta={\mathcal{P}}(\overline{I})$, where $\overline{I}$ is the complex conjugate interval, and $\theta$ commutes with $V_{t}$.
@@
[…]
@@
The function $G(\tau)=F_{2}(-\tau+i\pi)$ is defined on the strip $\{\pi\leq\operatorname{Im}(\tau)\leq 2\pi\}$, and $G(t+i\pi)=F_{1}(t+i\pi)$ for all $t\in{\mathbb{R}}$. Hence $F_{1}$ and $G$ assemble to a function $F$ continuous on ${\mathbb{S}}_{2\pi i}$ and holomorphic in the interior such that
@@
The challenges described above can make it difficult to study duality phenomena in the present context, which in turn increases the difficulty in precisely identifying the domains of various operators. For example, a key question that we do not resolve in this article is the following.
@@
We will see below (Corollary 5.13 and Proposition 5.18) that the answer to Question 4.4 has meaningful consequences for our present study.
@@
$G(t)=\langle V_{t}\Phi,\Phi^{\prime}\rangle,\qquad G(t+\tau_{0})=\langle V_{\tau_{0}}\Phi,V_{-t}\Phi^{\prime}\rangle.$
Similarly, since $\Phi^{\prime}\in D(V_{\tau_{0}})$ we have a map $H$ on the same strip (continuous on the closure and holomorphic on the interior) such that
@@
On the domain $D(V_{i\pi})$ introduced in Definition 4.1, $V_{i\pi}^{*}=V_{i\pi}$. Similarly, $V_{-i\pi}^{*}=V_{-i\pi}$.
@@
We only consider $V_{i\pi}$. Fix $\Phi^{\prime}\in D(V_{i\pi}^{*})$, and we must show that $\Phi^{\prime}\in D(V_{i\pi})$. Let $\Phi\in D(V_{i\pi})$, and let $G:{\mathbb{S}}\to{\mathbb{C}}$ be the continuous function (holomorphic in the interior) such that for all $t\in{\mathbb{R}}$ we have
@@
[…]
$G(t+i\pi)=\langle V_{t}V_{i\pi}\Phi,\Phi^{\prime}\rangle=\langle\Phi,V_{-t}V_{i\pi}^{*}\Phi^{\prime}\rangle.$
@@
$H_{\Phi}(t)=\langle V_{t}\Phi^{\prime},\Phi\rangle,\qquad H_{\Phi}(t+i\pi)=\langle V_{t}V_{i\pi}^{*}\Phi^{\prime},\Phi\rangle.$ (4.3)
In order to show that $\Phi^{\prime}\in D(V_{i\pi})$, we must establish the existence of such a function $H_{\Phi}$ for all $\Phi\in{\mathcal{D}}$, and we have done this so far when $\Phi\in D(V_{i\pi})$.
By Lemma 3.4, the closure of ${\mathcal{P}}_{\mathcal{X}}(I_{+})\Omega$ contains ${\mathcal{P}}(I_{+})\Omega$, and thus ${\mathcal{P}}_{\mathcal{X}}(I_{+})\Omega$ is dense by the Reeh-Schlieder theorem [CRTT24, App. A]. Given arbitrary $\Phi\in{\mathcal{D}}$, choose a net $\Phi_{n}\in{\mathcal{P}}_{\mathcal{X}}(I_{+})\Omega$ such that $\Phi_{n}\to\Phi$. We will show that the net $H_{\Phi_{n}}$ converges locally uniformly on ${\mathbb{S}}$ to the required function $H_{\Phi}$. Our first step is to show that this net is locally uniformly Cauchy.
Note that $\overline{H_{\Phi_{n}}(-\overline{\tau})}$ is a function of the form $G$ considered above, and so by Lemma 3.9 (and the fact that $\Phi_{n}\in{\mathcal{P}}_{\mathcal{X}}(I_{+})\Omega$) we can infer that $e^{-2\tau^{2}}H_{\Phi_{n}}$ vanishes at infinity. It follows that $e^{-2\tau^{2}}H_{\Phi_{n}}-e^{-2\tau^{2}}H_{\Phi_{m}}(\tau)$ vanishes at infinity as well, and so
$\left|e^{-2\tau^{2}}H_{\Phi_{n}}(\tau)-e^{-2\tau^{2}}H_{\Phi_{m}}(\tau)\right|$ (4.4)
is maximized on $\partial{\mathbb{S}}={\mathbb{R}}\cup i\pi+{\mathbb{R}}$. From (4.3), we see that the restriction of (4.4) to each boundary component is a function ${\mathbb{R}}\to{\mathbb{C}}$ of the form
$t\mapsto Ce^{-2t^{2}}\langle V_{t}\Psi,\Phi_{n}-\Phi_{m}\rangle$
for some $\Psi\in{\mathcal{D}}$. We will show that functions of the above form converge uniformly to zero as $n,m\to\infty$.
@@
From the proof of Lemma 3.8 (specifically (3.10)), we have $e^{-2t^{2}/k}\beta_{d}(v_{t})f\to 0$ in $C^{\infty}(S^{1})$ as $t\to\pm\infty$. Hence $e^{-2t^{2}}V_{t}\Psi\to 0$ in ${\mathcal{D}}$, or in other words $e^{-t^{2}}V_{t}\Psi$ extends to a continuous function $[-\infty,\infty]\to{\mathcal{D}}$ which vanishes at infinity. Since [,][$\infty$,$\infty$][ - ∞ , ∞ ] is compact, the image of this map is a bounded subset of ${\mathcal{D}}$, which means that for any open neighborhood $U$ of 00 in ${\mathcal{D}}$ there exists a scalar $M>0$ such that $e^{-t^{2}}V_{t}\Psi\in MU$ for all $t$.
Since the Hermitian form is jointly continuous, given $\epsilon>0$ we may choose a neighborhood $U$ of 00 in ${\mathcal{D}}$ such that $\left|\langle\Phi_{1},\Phi_{2}\rangle\right|<\epsilon$ when $\Phi_{j}\in U$. Choose $M>0$ as above so that $V_{t}\Psi\in MU$ for all $t$, and note that $\left|\langle\Phi_{1},\Phi_{2}\rangle\right|<\epsilon$ when $\Phi_{1}\in MU$ and $\Phi_{2}\in M^{-1}U$. Since $\Phi_{n}\to\Phi$, we may choose $n_{0}$ such that when $n,m\geq n_{0}$ we have $\Phi_{n}-\Phi_{m}\in M^{-1}U$. We then have for all $t\in{\mathbb{R}}$
@@
and so the sequence of functions $e^{-2t^{2}}\langle V_{t}\Psi,\Phi_{n}-\Phi_{m}\rangle$ converges uniformly to zero, as claimed. As described above, it follows that $e^{-\tau^{2}}(H_{\Phi_{n}}-H_{\Phi_{m}})$ converges uniformly to zero. Hence $H_{\Phi_{n}}$ is a locally uniformly Cauchy net of functions, which therefore converges locally uniformly to a function $H_{\Phi}$ that is continuous on ${\mathbb{S}}$ and holomorphic in the interior. Taking pointwise limits on the boundary we see that $H_{\Phi}$ satisfies (4.3). Since $\Phi$ was arbitrary, we conclude that $\Phi^{\prime}\in D(V_{i\pi})$, as claimed. ∎

### v2

@@ 3 Analytic continuation for non-unitary Wightman CFTs
We have thus shown that for $\tau\in{\mathbb{S}}$ the domain $D(\widetilde{V}_{\tau})$ contains ${\mathcal{P}}_{\mathcal{X}}(I_{+})\Omega$, where we recall ${\mathcal{P}}_{\mathcal{X}}(I_{+})\subset{\mathcal{P}}(I_{+})$ is generated by operators $\varphi(F|_{I_{+}})$ with $F\in{\mathcal{X}}$. We would like to extend this result to ${\mathcal{P}}(I_{+})\Omega$ via a limiting argument, approximating $f\in C^{\infty}_{0}(I_{+})$ by functions of the form $F|_{I_{+}}$. We will begin by examining the growth rate of the function $\lambda(V_{t}\Phi)$.
@@
    where $\left\|f\right\|_{C^{N}}=\sum_{j=0}^{N}\big\|\tfrac{d^{j}}{d\theta^{j}}f\big\|_{\infty}$.
@@
is jointly continuous on $C^{\infty}(S^{1})^{k}$. Thus by standard results (cf. the beginning of [Trè67, §41]) there exists a positive integer $N$ such that
@@
where we take the convention that the (non-commuting) products are ordered left-to-right, and the empty products (when $j=1$ or $j=k$) are omitted. Hence
@@
$\qquad\leq\sum_{j=1}^{k}\left|\lambda\left(\left[\prod_{\ell=1}^{j-1}\varphi_{\ell}(\beta_{d_{\ell}}(v_{t})f_{\ell})\right]\varphi_{j}(\beta_{d_{j}}(v_{t})[f_{j}-g_{j}])\left[\prod_{\ell=j+1}^{k}\varphi_{\ell}(\beta_{d_{\ell}}(v_{t})g_{\ell})\right]\Omega\right)\right|\leq$
@@
by Lemma 3.7(i), and similarly for the $g_{j}$. By Lemma 3.8 applied to the functions $f_{j,s}$ and $g_{j,s}$ (along with the fact that $\widetilde{V}_{t+is}=V_{t}\widetilde{V}_{is}$) we have
@@
for some positive integer $N$ and some function $C$ with at most exponential growth that does not depend on the functions $F_{j}$ and $G_{j}$. Since $F_{j}$ and $G_{j}$ are smooth on $\mathbb{O}$}, the maximum
@@
is achieved at some $s_{0}\in[0,\pi]$. Hence there is a constant $K>0$ (depending on the $F_{j}$ and $G_{j}$) such that
@@
where $M=\max_{t}e^{-t^{2}}C(t)$ does not depend on the functions $F_{j}$$G_{j}$.
@@
$\big|\lambda\big(\big[\widetilde{V}_{\tau}$ $\varphi_{1}(f_{1,n})\cdots\varphi_{k}(f_{k,n})\Omega\big]-\big[\widetilde{V}_{\tau}\varphi_{1}(f_{1,m})\cdots\varphi_{k}(f_{k,m})\Omega\big]\big)\big|$
@@
$G_{\lambda,n}(\tau)=\lambda\big(\widetilde{V}_{\tau}\varphi_{1}(f_{1,n})\cdots\varphi_{k}(f_{k,n})\Omega\big)$
@@ 4 Invariant Hermitian forms, Bisognano-Wichmann, and the KMS condition
In this section, we assume that our Wightman CFT $({\mathcal{F}},{\mathcal{D}},U,\Omega)$ has an invariant nondegenerate sesquilinear form $\langle\,\cdot\,$$\cdot\,\rangle$, along with an antiunitary PCT involution $\theta:{\mathcal{D}}\to{\mathcal{D}}$ which fixes $\Omega$. Such a theory is called involutive (see [CRTT25, §4] for additional background and the connection with vertex algebras). The invariance of the form $\langle\,\cdot\,$$\cdot\,\rangle$ means, firstly, that
@@
and moreover we have $(-1)^{d_{\varphi}}\varphi^{\dagger}\in{\mathcal{F}}$. Secondly, an invariant form must also satisfy
$\langle U(\gamma)\Phi,U(\gamma)\Psi\rangle=\langle\Phi,\Psi\rangle,\qquad\Phi,\Psi\in{\mathcal{D}},\,\gamma\in\operatorname{M\ddot{o}b}.$ (4.2)
The sesquilinear form is assumed to be jointly $\mathcal{F}$}-strong continuous and is automatically Hermitian symmetric, meaning $\langle\Phi,\Psi\rangle=\overline{\langle\Psi,\Phi\rangle}$ for all $\Phi,\Psi\in{\mathcal{D}}$. Given $\theta$, the existence of a sesquilinear form on $\mathcal{D}$} with these properties is equivalent to $\langle\,\cdot\,,\theta\,\cdot\,\rangle$ being a nondegenerate invariant bilinear form on the Möbius vertex algebra underlying $\mathcal{F}$}. The theory is called unitary if the Hermitian form is an inner product (normalized so that $\left\|\Omega\right\|=1$). In terms of the correspondence between Wightman CFTs and vertex algebras, an involutive Wightman CFT corresponds to a Möbius vertex algebra with a sesquilinear form and antiunitary operator $\theta$ such that $\langle\,\cdot\,\theta\,\cdot\,\rangle$ is an invariant bilinear form in the sense of [FHL93, Li94]. See [CRTT25, Def. 4.9] and the surrounding discussion for more detail.
Recall that $\operatorname{End}({\mathcal{D}})$ denotes the collection of $\mathcal{F}$}-strong continuous linear operators on $\mathcal{D}$}. Given $x\in\operatorname{End}({\mathcal{D}})$, we say that $x$ has an adjoint if there exists $x^{*}\in\operatorname{End}({\mathcal{D}})$ such that $\langle x\Phi,\Psi\rangle=\langle\Phi,x^{*}\Psi\rangle$ for all $\Phi,\Psi\in{\mathcal{D}}$ [^8] See also the discussion following Corollary 4.3 regarding the partially defined adjoint of a partially defined (and not necessarily continuous) linear map., and write $\operatorname{End}_{*}({\mathcal{D}})$ for the collection of operators $x\in\operatorname{End}({\mathcal{D}})$ that have an adjoint $x^{*}\in\operatorname{End}({\mathcal{D}})$. In light of (4.1) we see that $\varphi(f)\in\operatorname{End}_{*}({\mathcal{D}})$, and moreover the algebras ${\mathcal{P}}(I)$ are *-algebras, i.e., they are closed under taking adjoints. We have $\theta{\mathcal{P}}(I)\theta={\mathcal{P}}(\overline{I})$, where $\overline{I}$ is the complex conjugate interval, and $\theta$ commutes with $V_{t}$. If $\gamma\in\operatorname{M\ddot{o}b}$, then (4.2) implies that $U(\gamma)\in\operatorname{End}_{*}({\mathcal{D}})$, with $U(\gamma)^{*}=U(\gamma^{-1})$.
@@
and let $\alpha_{t}=\operatorname{Ad}V_{t}\in\operatorname{Aut}({\mathcal{P}}(I_{+}))$. Let ${\mathbb{S}}_{2\pi i}=\{\tau\in{\mathbb{C}}\,:\,0\leq\operatorname{Im}(\tau)\leq 2\pi\}$. Then for every $x,y\in{\mathcal{P}}(I_{+})$, there exists a function $F:{\mathbb{S}}_{2\pi i}\to{\mathbb{C}}$ that is continuous on ${\mathbb{S}}_{2\pi i}$, holomorphic on $\mathring{{\mathbb{S}}}_{2\pi i}$, and satisfies for all $t\in{\mathbb{R}}$
@@
Let $G:\{\pi\leq\operatorname{Im}(\tau)\leq 2\pi\}\to{\mathbb{C}}$ be the function $G(\tau)=F_{2}(-\tau+i\pi)$. We can compute
$G(t+i\pi)=F_{2}(-t)=\langle V_{-t}\theta x^{*}\Omega,y^{*}\Omega\rangle=\langle\theta V_{t}y^{*}\Omega,x^{*}\Omega\rangle=F_{1}(t+i\pi)$
for all $t\in{\mathbb{R}}$, where we have used the fact that $V_{t}^{*}=V_{-t}$ in the third equality and the fact that $\theta$ and $V_{t}$ commute in the final equality. Hence $F_{1}$ and $G$ assemble to a function $F$ continuous on ${\mathbb{S}}_{2\pi i}$ and holomorphic in the interior such that
@@
The challenges described above can make it difficult to study duality phenomena in the present context, which in turn increases the difficulty in precisely identifying the domains of various operators. For example, a question that we do not resolve in this article is the following.
@@
$G(t)=\langle V_{t}\Phi,\Phi^{\prime}\rangle,\qquad G(t+\tau_{0})=\langle V_{t}V_{\tau_{0}}\Phi,\Phi^{\prime}\rangle=\langle V_{\tau_{0}}\Phi,V_{-t}\Phi^{\prime}\rangle,$
where the last equality follows from (4.2). Similarly, since $\Phi^{\prime}\in D(V_{-\overline{\tau_{0}}})$ we have a map $H$ on the same strip (continuous on the closure and holomorphic on the interior) such that
@@
Let $({\mathcal{F}},{\mathcal{D}},U,\Omega)$ be an involutive Wightman CFT on $S^{1}$, and let $V_{\pm i\pi}$ be the densely defined operators from Definition 4.1 with domains $D(V_{\pm i\pi})$. Then we have $V_{\pm i\pi}^{*}=V_{\pm i\pi}$. Moreover, the adjoints of the restrictions $V_{\pm i\pi}|_{{\mathcal{P}}(I_{\pm})\Omega}$ are again given by $\big(V_{\pm i\pi}|_{{\mathcal{P}}(I_{\pm})\Omega}\big)^{*}=V_{\pm i\pi}$, where ${\mathcal{P}}(I_{\pm})$ are the unital algebras generated by operators $\varphi(f)$ with $\varphi\in{\mathcal{F}}$ and $\operatorname{supp}f\subset I_{\pm}$.
@@
We only consider $V_{i\pi}$. Observe that we have a chain of extensions of densely defined operators
$V_{i\pi}|_{{\mathcal{P}}(I_{+})\Omega}\subseteq V_{i\pi}\subseteq V_{i\pi}^{*}\subseteq(V_{i\pi}|_{{\mathcal{P}}(I_{+})\Omega})^{*},$
where the first inclusion is immediate, the second inclusion is Lemma 4.5, and the third follows by taking adjoints in the first inclusion. To prove the theorem, it therefore suffices to show that $(V_{i\pi}|_{{\mathcal{P}}(I_{+})\Omega})^{*}\subset V_{i\pi}$. In fact, since we know that these two operators agree on $D(V_{i\pi})$ we need only show the corresponding inclusion of domains:
$D\big((V_{i\pi}|_{{\mathcal{P}}(I_{+})\Omega})^{*}\big)\subseteq D(V_{i\pi}).$
Fix $\Phi^{\prime}\in D\big((V_{i\pi}|_{{\mathcal{P}}(I_{+})\Omega})^{*}\big)$, and we must show that $\Phi^{\prime}\in D(V_{i\pi})$. Let $\Phi\in{\mathcal{P}}(I_{+})\Omega$. Since ${\mathcal{P}}(I_{+})\Omega\subset D(V_{i\pi})$, there exists a function $G:{\mathbb{S}}\to{\mathbb{C}}$ which is continuous, holomorphic in the interior $\mathring{{\mathbb{S}}}$, and satisfies for all $t\in{\mathbb{R}}$
@@
where as usual ${\mathbb{S}}=\{\tau\in{\mathbb{C}}\,:\,0\leq\operatorname{Im}(\tau)\leq\pi\}$. Let $\Phi^{\prime\prime}=(V_{i\pi}|_{{\mathcal{P}}(I_{+})\Omega})^{*}\Phi^{\prime}$. Note that $V_{t}$ leaves ${\mathcal{P}}(I_{+})\Omega$ invariant and commutes with $V_{i\pi}$, so for $t\in{\mathbb{R}}$ we have
$G(t+i\pi)=\langle V_{i\pi}V_{t}\Phi,\Phi^{\prime}\rangle=\langle V_{t}\Phi,\Phi^{\prime\prime}\rangle=\langle\Phi,V_{-t}\Phi^{\prime\prime}\rangle.$
@@
$H_{\Phi}(t)=\langle V_{t}\Phi^{\prime},\Phi\rangle,\qquad H_{\Phi}(t+i\pi)=\langle V_{t}\Phi^{\prime\prime},\Phi\rangle.$ (4.4)
In order to show that $\Phi^{\prime}\in D(V_{i\pi})$, we must establish the existence of such a function $H_{\Phi}$ for all $\Phi\in{\mathcal{D}}$, and we have done this so far when $\Phi\in{\mathcal{P}}(I_{+})\Omega$.
By Lemma 3.4, the closure of ${\mathcal{P}}_{\mathcal{X}}(I_{+})\Omega$ contains ${\mathcal{P}}(I_{+})\Omega$, and thus ${\mathcal{P}}_{\mathcal{X}}(I_{+})\Omega$ is dense in $\mathcal{D}$} by the Reeh-Schlieder theorem [CRTT25, App. A]. Given arbitrary $\Phi\in{\mathcal{D}}$, choose a net $\Phi_{n}\in{\mathcal{P}}_{\mathcal{X}}(I_{+})\Omega$ such that $\Phi_{n}\to\Phi$. We will show that the net $H_{\Phi_{n}}$ converges locally uniformly on $\mathbb{S}$} to the required function $H_{\Phi}$. Our first step is to show that this net is locally uniformly Cauchy.
Note that $\overline{H_{\Phi_{n}}(-\overline{\tau})}$ is a function of the form $G$ considered above, and so by Lemma 3.9 (and the fact that $\Phi_{n}\in{\mathcal{P}}_{\mathcal{X}}(I_{+})\Omega$) we can infer that $e^{-2\tau^{2}}H_{\Phi_{n}}$ vanishes at infinity. It follows that $e^{-2\tau^{2}}H_{\Phi_{n}}(\tau)-e^{-2\tau^{2}}H_{\Phi_{m}}(\tau)$ vanishes at infinity as well, and so the magnitude of
$e^{-2\tau^{2}}H_{\Phi_{n}}(\tau)-e^{-2\tau^{2}}H_{\Phi_{m}}(\tau)$ (4.5)
is maximized on $\partial{\mathbb{S}}={\mathbb{R}}\cup i\pi+{\mathbb{R}}$. From (4.4), we see that the restrictions of (4.5) to the components of $\partial{\mathbb{S}}$ are given by
$t\mapsto e^{-2t^{2}}\langle V_{t}\Phi^{\prime},\Phi_{n}-\Phi_{m}\rangle,\qquad\mbox{and}\qquad t+i\pi\mapsto e^{2\pi^{2}-4\pi it}e^{-2t^{2}}\langle V_{t}\Phi^{\prime\prime},\Phi_{n}-\Phi_{m}\rangle,$
respectively. Thus to show that (4.5) converges uniformly to zero on $\mathbb{S}$}, it suffices to show that for any $\Psi\in{\mathcal{D}}$, functions of the form
$t\mapsto e^{-2t^{2}}\langle V_{t}\Psi,\Phi_{n}-\Phi_{m}\rangle$
converge uniformly to zero on $\mathbb{R}$} as $n,m\to\infty$.
@@
From the proof of Lemma 3.8 (specifically (3.10)), we have $e^{-2t^{2}/k}\beta_{d}(v_{t})f\to 0$ in $C^{\infty}(S^{1})$ as $t\to\pm\infty$. Hence $e^{-2t^{2}}V_{t}\Psi\to 0$ in $\mathcal{D}$}, or in other words $e^{-2t^{2}}V_{t}\Psi$ extends to a continuous function $\infty,\infty]\to{\mathcal{D}}$ which vanishes at infinity. Since $\infty$$\infty$] is compact, the image of this map is a bounded subset of $\mathcal{D}$}, which means that for any open neighborhood $U$ of 0 in $\mathcal{D}$} there exists a scalar $M>0$ such that $e^{-2t^{2}}V_{t}\Psi\in MU$ for all $t$.
Since the Hermitian form is jointly continuous, given $\epsilon>0$ we may choose a neighborhood $U_{\epsilon}$ of 0 in $\mathcal{D}$} such that $\left|\langle\Phi_{1},\Phi_{2}\rangle\right|<\epsilon$ when $\Phi_{j}\in U_{\epsilon}$. Choose $M>0$ as above so that $e^{-2t^{2}}V_{t}\Psi\in MU_{\epsilon}$ for all $t$, and note that $\left|\langle\Phi_{1},\Phi_{2}\rangle\right|<\epsilon$ when $\Phi_{1}\in MU_{\epsilon}$ and $\Phi_{2}\in M^{-1}U_{\epsilon}$. Since $\Phi_{n}\to\Phi$, we may choose $n_{0}$ such that when $n,m\geq n_{0}$ we have $\Phi_{n}-\Phi_{m}\in M^{-1}U_{\epsilon}$. We then have for all $t\in{\mathbb{R}}$
@@
and so the net of functions $e^{-2t^{2}}\langle V_{t}\Psi,\Phi_{n}-\Phi_{m}\rangle$ converges uniformly to zero, as claimed. As described above, it follows that $e^{-2\tau^{2}}(H_{\Phi_{n}}-H_{\Phi_{m}})$ converges uniformly to zero. Hence $H_{\Phi_{n}}$ is a locally uniformly Cauchy net of functions, which therefore converges locally uniformly to a function $H_{\Phi}$ that is continuous on $\mathbb{S}$} and holomorphic in the interior. Taking pointwise limits on the boundary we see that $H_{\Phi}$ satisfies (4.4). Since $\Phi$ was arbitrary, we conclude that $\Phi^{\prime}\in D(V_{i\pi})$, as claimed. ∎
As a consequence of Theorem 4.6, we have $\big(V_{\pm i\pi}|_{{\mathcal{P}}(I_{\pm})\Omega}\big)^{**}=V_{\pm i\pi}$. In the context of Hilbert spaces, the double adjoint coincides with the closure, and so the ‘moreover’ of the Theorem may be thought of as a positive answer to a weakening of Question 4.4.